PREX/CREX Oct 16, 2018 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  555 144 462
 Bluejeans link:
 Offline room:   CEBAF Center F324-325


  1. analyzer update: Ciprian (et al.)
    • CG: working with Ye and Musson to figure out why we can't readout the BPM signals properly. We suspect a problem with the DAC configurations.
  2. heating and radiation around collimator: Adam HAPLOG Entry 3552
    • AZ: we are looking at the dose first. We had determined that the dose on the plastic shielding around the collimator was very high. Now we cut out a piece of the plastic (from 4.25 in to 6 in) and also went to flat momentum raster (momentum along Z only). Looked at the total dose on the shield and there was a drop (10%). KP: the circumference got significantly larger but it’s not as significant as we had hoped.
    • KP: the power in the flange and cone DS of the collimator box is slightly smaller because we had a raster that was not quite right. Talked to Robin about it and she thinks it’s not a huge amount to of heating to remove by bringing a cooling tube in. Not clear if we need any copper on top. I am leaving it to the engineers to figure out the details.
  3. New collimator design haplog3536, haplog3553
    • KP: the sieves was put wrong in the model and we have a new design that has the proper position. In particular the sieve was not perpendicular to the acceptance envelope.
    • KP: we have transmitted the new sieve design to the engineering team. We decided to remove the top row and right column since they were outside of the acceptance.
    • KP: we are making sure that we have the center position correct since Adam and Seamus have different results (to a couple of mm).
    • SR: let’s take this offline and make sure the calculations are correct. It’s a geometry problem and we should get the same result for PREX and CREX. DocDB 181 has the FOM calculation and should be a good indicator of I calculated.
  4. SAMs: Dustin
    • DM: could we have a SAM meeting this Th? Looking to finalize the light guide part of the construction and prepare for a test beam. SR: I’ll arrange it.
  5. GEMs
    • KP: how are the GEM prep going? DM: everything is progressing well and we are systematically developing software tools for it. We are investigating some issues with our source. Our cosmics are going well. Seamus is working on the gas side of things and things are going well.
    • DM: we’ll still have to figure out the DAQ issues, but we’ll have a meeting on this before Dec.


Amali P, Sanghwa P, Dave G, Ciprian G, Victoria O, Adam Z, Chandan G, David A, Devi A, Dustin M, Guido U, KK, Rich Holmes, Ryan R, Bob M, Seamus R, Silviu D, Tao Y, Kent P, Ye T, Cameron C, Eric K, Jinlong Z, Sakib R, Siyu Jian, Caryn P,